OpenStudy (anonymous):

Find the indefinite integral. \int [x^{2}+\frac{1}{3x^2}]dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\left[ x^2+\left(\begin{matrix}1 \\ 3x^2\end{matrix}\right) \right]dx\]

OpenStudy (anonymous):

\[=\int\limits x^2~dx+\frac{ 1 }{ 3 }\int\limits x ^{-2}dx+c\] \[\int\limits x^n dx=\frac{ x ^{n+1} }{ n+1 }\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw